STOCK VALUATION:

Price–Earnings Ratio Consider Pacific Energy Company and U.S. Bluechips, Inc., both of which reported earnings of $630,000. Without new projects, both firms will continue to generate earnings of $630,000 in perpetuity. Assume that all earnings are paid as dividends and that both firms require a return of 11 percent. a. What is the current PE ratio for each company? b. Pacific Energy Company has a new project that will generate additional earnings of $100,000 each year in perpetuity. Calculate the new PE ratio of the company. c. U.S. Bluechips has a new project that will increase earnings by $200,000 in perpetuity. Calculate the new PE ratio of the firm.
Answer #1

Earnings 
630000GIVEN











Perpetual Earnings 
630000











RATE OF Return
11%


























A)FORMULA FOR P/E RATIO =   PRICE / EARNINGS











WE KNOW ONLY EARNINGS WE   DON’T NO THE PRICE SO 1ST WE HAVE TO CALCULATE THAT.






FORMULA FOR PRICE =   C/(r-g)
HERE   "C" IS DIVIDEND... AS IT IS SAID IN QUESTION THAT ALL THE EARNINGS   ARE GIVEN AS DIVIDEND… SO, WE WILL TAKE EARNINGS AS DIVIDEND. AND   "GROWTH" IS NOT GIVEN SO WE WILL TAKE IT AS 0 


PRICE =5727272.73


PE RATIO =9.09



























B)NEW EARNINGS = 100000












 TOTAL EARNINGS = 100000+630000730000NOW THEY ARE GIVING THIS AS DIVIDEND






















FORMULA FOR P/E RATIO =   PRICE / EARNINGS











WE KNOW ONLY EARNINGS WE   DON’T NO THE PRICE SO 1ST WE HAVE TO CALCULATE THAT.






FORMULA FOR PRICE =   C/(r-g)













PRICE =6636363.64













PE RATIO =10.53



























C)NEW EARNINGS = 200000












 TOTAL EARNINGS= 200000+630000830000NOW THEY ARE GIVING THIS AS DIVIDEND






















FORMULA FOR P/E RATIO =   PRICE / EARNINGS











WE KNOW ONLY EARNINGS WE   DON’T NO THE PRICE SO 1ST WE HAVE TO CALCULATE THAT.






FORMULA FOR PRICE =   C/(r-g)





























PRICE =7545454.55













PE RATIO =11.98




























answered by: Sai Nath
Answer #2

SOLUTION :


Current value of both companies 

= PV of future cash flows of $63000 in perpetuity discounted at 11%.

= 630000 / 0.11

= 5727272.72 ($)


Let number of shares be N.


So, 


Price per share = 5727272.72 / N  ($)


Earring per share = 630000 / N   ($)


PE ratio of each company.

= (5727272.72 / N) / (630000 / N) 

= 5727272.72 / 630000 

= 9.09 (ANSWER). 


With new project, earnings of Pacific energy company = 630000 + 100000 = 730000 


PE ratio of this company 

= Value of company / Earnings

= (730000 / 0.11) / 730000 

= 9.09  (ANSWER) 


With new project, earnings of US bluechip company = 630000 + 200000 = 830000 


PE ratio of this company 

= Value of company / Earnings

= (830000 / 0.11) / 830000 

= 9.09  (ANSWER) 

answered by: Tulsiram Garg

> Please correct in the 2nd line , "63000" as "630000" . Thanks.

Tulsiram Garg Fri, Oct 29, 2021 8:26 AM